OpenStudy (anonymous):

pleas find the Laplace transform this function t^2e^tsin3t

OpenStudy (experimentx):

\[ L(t^n f(t)) = (-1)^n F^n(s) \] http://www.wolframalpha.com/input/?i=Laplace+Transform+t^2+e^t+sint

OpenStudy (anonymous):

oh thanks for u dear its good solutions thanks

OpenStudy (experimentx):

not a solution .. just a practice for me

OpenStudy (experimentx):

\[ \large L(\sin 3t e^t) = \frac 3 {(s-1)^2+3^2}\]

OpenStudy (anonymous):

L(t^n f(t)) = (-1)^n dn/ds^n F(s) this equation its coplex differentiation but i dont solution tis problem but f(s)=?

OpenStudy (experimentx):

\[ \huge F(s) = \large L(\sin 3t e^t) = \frac 3 {(s-1)^2+3^2} \]

OpenStudy (anonymous):

thanks for u dear i solve this thanks

OpenStudy (experimentx):

you are welcome !!